# EVEv2: Improved Baselines for Encoder-Free Vision-Language Models

### v1

@@ Appendix A Experiment Details
For Exp.(ii) in Figure 2, all VLMs use stronger Qwen2-7B [89] and high-quality EVE-recap. In Stage 1, we train the projector for VE, patch embedding layer for EVEv1.0, patch embedding and extra vision layer inside the LLM for EVEv1.2. In Stage 2 and 3, we train all model weights for EVEv1.0-1.2. Note that we skip Stage 2 for VE as the baseline for comparison. Here, we compare the weights between Qwen2 and VE/EVEv1.2 trained by LLaVA-onevision [38].
Appendix B Dataset Details
The dataset details in Stages 2.2 and 3 are listed in Table 4. In Stage 2.2, we adopt Cambrian-FL [71], Infinity-Instruct-FL [27], LVIS-Instruct-FL [77], Sharegpt4v-FL [12], ALLaVA-laion-FL [11], ALLaVA-vflan-FL [11], LLaVA-Pretrain-FL [47], DocReason-FL [29], DocDownstream-FL [29], and DocStruct4M-FL [29]. In Stage 3, we adopt LLaVA-onevision [38], Infinity-MM-Synthesis [27], Infinity-MM-Preference [27], Infinity-Instruct-FL [27], DenseFusion [43], Cambrian-FL [71], Docmatix-FL [36], LVIS-Instruct-FL [77], BLIP-OCR[17], and LLaVA-mix [47].
Appendix C Hyper-parameter Configurations
Appendix D Visual Understanding Demonstration
We investigate several vision perception and reasoning capabilities of EVEv2.0, including OCR capability in Tables 6 and 7, mixed information in Table 8, real-world scenes in Tables 9 and 10, and set-of-mark prompting task in Table 11.
@@ Appendix D Visual Understanding Demonstration
Example 2: Understanding OCR Information in PPT [Uncaptioned image] User Please output all the text information. EVEv2.0 (A) Modular Large Vision-Language Model (B) Originated Large Vision-Language Model Language Model Interaction Following (i) Aligning Reasoning (ii) Differentiating Vision Encoder Perceiver Merging Projector Abstractor Vision Encoder Vision Encoder Existing Issues (1) Resolution/Aspect Ratio (2) Deployment Overhead (3) Capacity Proportion Motivation: Can We Train an Encoder-free LVLM? How can We Efficiently Build Originated LVLMs that can Rival Modular LVLMs? Property: (1) Decoder-only Architecture (2) Simpler LVLM to Scale & Deploy EVE.

### v2

@@ Appendix C Hyper-parameter Configurations
Appendix D Additional Ablation Studies
Decreasing SQA-Img metric with more training data. We empirically discovered that the performance of the SQA-Img metric is strongly influenced by the knowledge of the pre-trained LLM, making it prone to catastrophic linguistic forgetting issues. As it primarily involves text-centric knowledge tasks, increasing image captioning data alone degrades performance when the LLM parameters are unfrozen.
EVEv1.2: lowest loss, poor performance in Figure 5. This is because EVEv1.2 progressively updates the LLM weights while quickly adapting to the simplified language structures within the VL caption data. Notably, the SFT losses—which better reflect task performance—exhibit a clear downward trend (EVEv1.2: 1.0306, EVEv1.5: 0.9513, EVEv2: 0.9327), indicating that EVEv1.2 is less effective at alleviating LLM forgetting and multimodal interference.
Consistent benefits across different LLM capacities. We train EVEv1, v1.5, and v2 using Qwen2.5-1.5B with EVE-recap-4M and LLaVA-mix-665k, achieving average scores of 35.1%, 37.3%, and 38.5% across six diverse vision-language benchmarks. This further confirms that sparse design consistently benefits models of different scales.
Table 6: Inference speed comparison of various EVE counterparts
Model EVEv1.0 EVEv1.2 EVEv1.5 EVEv2.0
TTFT (s) 3.77 3.77 4.01 4.25
Inference speed comparisons across EVE variants. EVEv2.0 decouples weights across modalities at each layer, totaling 2$\times$7B parameters with active FLOPs per token matching a 7B dense model. Notably, we adopt an NVIDIA A100 with PyTorch, fixing image, input text, and output text tokens to 2500, 200, and 100, respectively. We report TTFT (time to first token) and TPS (tokens per second after the first token). Here, EVEv1.5 / EVEv2 bring marginal 6 / 13% TTFT overheads, while all EVEv1–v2 models sustain 35 TPS (tok/s) using KV-cache for visual and text prefixes.
Appendix E Visual Understanding Demonstration
We investigate several vision perception and reasoning capabilities of EVEv2.0, including OCR capability in Tables 7 and 8, mixed information in Table 9, real-world scenes in Tables 10 and 11, and set-of-mark prompting task in Table 12.
@@ Appendix E Visual Understanding Demonstration
Example 2: Understanding OCR Information in PPT
User Please output all the text information.
EVEv2.0 (A) Modular Large Vision-Language Model (B) Originated Large Vision-Language Model Language Model Interaction Following (i) Aligning Reasoning (ii) Differentiating Vision Encoder Perceiver Merging Projector Abstractor Vision Encoder Vision Encoder Existing Issues (1) Resolution/Aspect Ratio (2) Deployment Overhead (3) Capacity Proportion Motivation: Can We Train an Encoder-free LVLM? How can We Efficiently Build Originated LVLMs that can Rival Modular LVLMs? Property: (1) Decoder-only Architecture (2) Simpler LVLM to Scale & Deploy EVE.